Mrs. Jones invested a total of $12,000, part at 10% and the rest at 5%. If the income from her 10% investment was 6 times as great as the income from 5% investment, how much was invested at each rate?
Answer by scott8148(6628)

let x="amount @ 10%", so 12000-x="amount @ 5%"

"the income from her 10% investment was 6 times as great as the income from 5% investment"
__ x(10%)=6(12000-x)(5%)

multiplying by 100 to clear percentages __ 10x=360000-30x __ 40x=360000 __ x=9000